OpenStudy (anonymous):

While crossing the Atlantic, sailors spot two mermaids 120° apart on each end of an island that is 6 miles away. How far apart are the mermaids around the outer edge of the island to the nearest tenth of a mile?

OpenStudy (jziggy):

 The circumference of the full circle is 2(6)pi which is 12pi. You just want a section of this circumference (120/360) = (1/3) of the total circumference. So... 12pi(1/3) = (12/3)pi = 4pi miles